A light plane must reach a speed of 33 m/s for takeoff. How long a runway is needed if the acceleration is 3.0 m/s2?
1
Expert's answer
2020-07-06T15:13:12-0400

By definition of the acceleration, to accelerate from the rest to speed "v=33m\/s" with the acceleration "a = 3m\/s^2" the plane needs the following amount of time:


"t = \dfrac{v}{a}"

During this time, plane covers the following distance, according to the kinematic law:


"d = \dfrac{at^2}{2}"

Substituting the expression for "t" into the last formula, get:


"d = \dfrac{v^2}{2a}"

Calculating the numerical value, obtain:


"d = \dfrac{33^2}{2\cdot3} = 181.5m"

Answer. The runway is 181.5 m.